FRS Home Page
-
–
S4
–
Pictures
–
Drawings
–
Cable Connections – TPC Electronics – Memo for preparation before beam time
Official Manual of TPC (
PDF )
Recent Log and Memo about TPC (2019) (External link)
Present Configuration (as of 2020/Oct/22)
- 2nd TPC at S4, just behind 2nd TUM-MUSIC.
- Gap (vertical) = 6 cm (This means -2400 V for the drift field voltage)
- Installed in the normal orientation.
- Calibration positions should be x = -12, 0, +12 mm and y = -6, 0, +6 mm (based on notes from 2016 beamtime)
- Maximum allowed HV: -2400 V (drift), +1100 V (anode), -700 V (calibration, PMT).
- For heavy ions, anode and PMT voltages should be much lower than the maximum values.
- picture
Status during previous beam time
- HV(anode) = +1050V, HV(drift) = -2400V, HV(grid-pmt) = -700V
- Analog raw signal around 150--200 mV. Vth(tpc module) around 50 mV
- Calibration file is /d/land4/sec/s533/tpc42_calib_0028.lmd (Logbook p41-44)
- Parameter sets used during experiment is found
here.
2020/June/05--June/08 (proton beam, WASA000)
- Applied voltages: anode HV = +1150 V, drift = -2400 V, grid-PMT = -xxx V
- Calibration: wasa2020june_tpc42calib_0022.lmd
- Amplifier has been changed to the high gain one.
2020/May/xx--May/xx (Pb beam, S469, very low energy)
- Applied voltages: anode HV = +dddd V, drift = -2400 V, grid-PMT = dddd V
- Calibration: dddd
- During S469 run, TPC42 was placed just behind TPC41, see log 2020/Apr/28.
2020/Apr/08--Apr/27 (Pb beam, S468, S482, S481)
- Applied voltages: anode HV = + 580V, drift = -2400 V, grid-PMT = -400 V.
- Calibration ....
2020/Feb/28--Mar/16 (Xe beam, S474, S459, S480)
- Applied voltages: anode HV = +700 V, drift = -2400 V, grid-PMT = -400 V.
- Calibration
picture1,
picture2 and
parameters (with mask trigger) in run0133 (Feb.29)
- Y position is now caculated properly even with different trigger (time reference is subtracted).
2019/Dec/09--Dec/19 (U beam)
- After changing primary beam from 40Ar to 238U.
- Applied voltages: anode HV = +550 V, drift = -2400V
2019/Nov/18--Dec/06 (Ar beam)
- Installed at S4 as TPC42.
- All signals were checked and ok, calibration scintllator was fine.
- Gap is 6cm, order: TPC->Fiber
- Applied voltages: anode HV = +800 V, drift = -2400V, fiber PMT = -500V.
2019/April/1--3 (107Ag, 290MeV/u, primary beam)
- Installed at S4 as TPC42. (normal orientation)
- All signals were checked and ok.
- Signal from the calibration scintillator was ok as well.
- Gap = 6cm (drift voltage = -2400V)
- Applied HV: -2400 V (drift), +700 V (anode), -(400--500) V (calibration PMT)
--
NisanYaprakBenal - 2026-10-26
Updated November 07, 2019
Messagelog
--
SagarRoy - 2021-06-26